# Simultaneously search for multi-target Galactic binary gravitational waves in reduced parameter space with LMPSO-CV

### v1

@@ III.2 $\mathcal{F}$-statistic degeneracy in the parameter space
represents the velocity of LISA relative to binary $A$, while $\lambda_{0}$ denotes LISA’s Ecliptic Longitude at the given time. Additionally, $\lambda_{A}$ represents the Ecliptic Longitude of binary $A$, and $\beta_{A}$ represents its Ecliptic Latitude. Assuming we use binary $A^{\prime}$ as the match waveform, due to the Doppler effect of LISA orbit motion, $f_{A^{\prime}_{do}}$ is its LISA response frequency
$f_{A^{\prime}_{do}}=f_{A^{\prime}}\frac{c+\nu_{A^{\prime}}}{c},$ (12)
@@
$\lambda_{A}^{\prime}$ and $\beta_{A}^{\prime}$ are the Ecliptic Longitude and Ecliptic Latitude of binary $A^{\prime}$, respectively. Making these two modulated frequencies equal ($f_{A_{do}}=f_{A^{\prime}_{do}}$), we can get the equation
@@
This equation pertains to the presence of a GW signal from binary $A$ in LISA data. When employing the $\mathcal{F}$-statistic to detect the signal, it describes the distribution of the $\mathcal{F}$-statistic in the parameter space. Considering the aforementioned equation, if the LISA data contains a signal with the frequency $f_{A}$ relative to the SSB, the maximum frequency $f_{A^{\prime}_{\text{max}}}$ in the $\mathcal{F}$-statistic degeneracy is
$f_{A^{\prime}_{max}}=f_{A}\frac{c+|\vec{\nu}|}{c-|\vec{\nu}|}\approx 1.0002\times f_{A}$ (16)
@@
when $\theta=\frac{\pi}{2}$, $\lambda_{A^{\prime}}-\lambda_{A}=\pi$, and $\beta_{A}=\beta_{A^{\prime}}=0$. Similarly, the minimum frequency $f_{A^{\prime}_{min}}$ in the $\mathcal{F}$-statistic degeneracy is
$f_{A^{\prime}_{min}}=f_{A}\frac{c-|\vec{\nu}|}{c+|\vec{\nu}|}\approx 0.9998\times f_{A}$ (17)
@@
when $\theta=-\frac{\pi}{2}$, $\lambda_{A^{\prime}}-\lambda_{A}=\pi$, and $\beta_{A}=\beta_{A^{\prime}}=0$. It determines the frequency range $[f_{A^{\prime}_{min}},\leavevmode\nobreak\ f_{A^{\prime}_{max}}]$ of the $\mathcal{F}$-statistic degeneracy generated by the GW signal which has frequency $f_{A}$ relative to the SSB. Eq. 16 and 17 show the frequency range of the $\mathcal{F}$-statistic degeneracy generated by the $f_{A}$ signal is proportional to $f_{A}$ itself. This results in higher $f_{A}$ exhibiting a more extensive distribution of the $\mathcal{F}$-statistic, while lower $f_{A}$ presents a narrower frequency distribution for the $\mathcal{F}$-statistic. We selected a set of parameters from the LDC1-4 injection sources catalog and assigned them to binary $A$, with the specific parameter values detailed in Tab. 2 (Appendix A). The $\mathcal{F}$-statistic degeneracy originating from binary $A$ is plotted in Fig. 4. In Fig. 4(a), $\beta_{A^{\prime}}=\beta_{A}$, and in Fig. 4(b), $\lambda_{A^{\prime}}=\lambda_{A}$. We utilized a spatial resolution of $\pi/40$ radians in the figures. $\theta$ represents the angle between LISA and binary $A$, providing a representation of LISA at various positions in its orbit. The $\mathcal{F}$-statistic degeneracy for binary $A$ spans a frequency range approximately from $0.9998\times f_{A}$ to $1.0002\times f_{A}$, corresponding to [$2.090406\times 10^{-3}$ Hz, $2.091243\times 10^{-3}$ Hz].
@@
Figure 4: The $\mathcal{F}$-statistic degeneracy from binary $A$ is derived according to Eq. 15. The parameters of binary $A$$f_{A},\leavevmode\nobreak\ \beta_{A},\leavevmode\nobreak\ \lambda_{A}$—are known, and the position of binary $A$ is denoted by red cross marks. In Fig. 4(a), $\beta_{A}^{\prime}=\beta_{A}$, and in Fig. 4(b), $\lambda_{A}^{\prime}=\lambda_{A}$. The horizontal axis represents the frequency ($f_{A^{\prime}}$) of the matched waveform, binary $A^{\prime}$, while the vertical axis represents its Ecliptic Longitude or Ecliptic Latitude ($\lambda_{A}^{\prime}$ or $\beta_{A}^{\prime}$). $\theta$ denotes the angle between LISA and binary $A$, with different $\theta$ values representing various positions of LISA relative to binary $A$.
@@ III.3.1 Individual signal
We generate a singular LISA response GW signal with the parameters of binary $A$ outlined in Tab. 2. Subsequently, we introduce instrument noise to acquire mock data and compute the $\mathcal{F}$-statistic degeneracy for data instances with diverse parameter positions. The intentional selection of binary $A$, distinguished by a high SNR, aims to minimize the influence of instrument noise. In Fig. 5, the distribution of degeneracy noise of binary $A$ across three parameter space ($f$, $\beta$, and $\lambda$) is illustrated. Each subfigure has a spatial resolution of $\pi/40$ radians. The colorbar indicates the $\mathcal{F}$-statistic value calculated at the specific parameter position. The red cross in the (2, 5) subfigure marks the real parameter position of binary $A$. The horizontal coordinate represents Ecliptic Longitude $\lambda$, while the vertical coordinate signifies Ecliptic Latitude $\beta$. The frequency interval between two adjacent subfigures corresponds to the minimum frequency resolution $\mathrm{d}f=1.59\times 10^{-8}$ Hz for the two-year detection. From the figure, it is evident that, in addition to the “primary peak” of the $\mathcal{F}$-statistic at the position of binary $A$, numerous “secondary peaks” with elevated $\mathcal{F}$-statistic values emerge. These secondary peaks represent the degeneracy noise mentioned earlier.
@@
Fig. 6(a) and Fig. 6(b) illustrate the distribution of the $\mathcal{F}$-statistic in two-dimensional parameter space. The frequency resolution is $\mathrm{d}f$, and the Ecliptic Latitude $\beta$ or the Ecliptic Longitude $\lambda$ resolution is $\pi/40$ radians. Parameters not displayed in the figures remain constant and correspond to binary $A$. The $\mathcal{F}$-statistic degeneracy in both figures (Fig. 6(a) and Fig. 6(b)) demonstrates excellent consistency with the theoretical predictions illustrated in Fig. 4(a) and Fig. 4(b).
@@
Figure 6: The red cross in the figures represent the position of binary $A$, while the colorbar indicate the $\mathcal{F}$-statistic value. The vertical coordinate corresponds to Ecliptic Longitude $\lambda$ or Ecliptic Latitude $\beta$, respectively. The $\mathcal{F}$-statistic degeneracy for binary $A$ spans a frequency range approximately from $0.9998\times f_{A}$ to $1.0002\times f_{A}$, corresponding to [$2.090406\times 10^{-3}$ Hz, $2.091243\times 10^{-3}$ Hz].
III.3.2 Two overlapped signals
@@
Signals that overlap in parameter space introduce complexity to the distribution of the $\mathcal{F}$-statistic. The proximity of two signals in parameter space can lead to interference, manifesting as a high $\mathcal{F}$-statistic peak at an incorrect parameter position, diminishing the value at the actual binary’s parameter position. In scenarios with substantial differences in SNR between the two binaries, the impact of the low SNR signal on the high SNR signal might not be conspicuous, and the low SNR signal could be discerned by subtracting the high SNR signal. However, when overlapped signals exhibit similar SNRs, their interference may yield a more pronounced $\mathcal{F}$-statistic peak at an incorrect parameter position. This complicates the iterative-subtraction strategy, as subtracting an incorrect waveform becomes a possibility. In our method, as long as the $\mathcal{F}$-statistic peak at the real binary’s parameter position does not disappear or deviate significantly, the removal of degeneracy noise, detailed in Sec. V, can effectively eliminate the erroneous $\mathcal{F}$-statistic peak. Building upon binary $A$, we increment its frequency by integer multiples of the minimum frequency resolution $\mathrm{d}f$ to create binary $B$ ($f_{B}=f_{A}+\Delta f=f_{A}+N\times\mathrm{d}f,\leavevmode\nobreak\ N\in Z$). Fig. 7 illustrates the distribution of the $\mathcal{F}$-statistic when two signals coexist. The frequency of binary $B$ in Fig. 7(a) and Fig. 7(b) is increased $5\times\mathrm{d}f$ relative to binary $A$, with red crosses denoting the accurate positions of both binaries—binary $A$ on the left and binary $B$ on the right. When the frequency difference $\Delta f$ between two signals is $\geq 5\times\mathrm{d}f$, the “primary peaks” of the $\mathcal{F}$-statistic for these two signals do not significantly influence each other, each behaving akin to a single source. However, for the binary $B$ in Fig. 7(c) and Fig. 7(d), where the frequency of binary $B$ is incremented by $4\times\mathrm{d}f$ relative to binary $A$, the overlap of signals causes a noticeable deviation in their $\mathcal{F}$-statistic behavior compared to the single-signal scenario. In Fig. 7(c), two maximal $\mathcal{F}$-statistic peaks emerge outside the parameter positions of binaries $A$ and $B$, substantially reducing the $\mathcal{F}$-statistic values at the true binary positions. Similarly, in Fig. 7(d), the maximal $\mathcal{F}$-statistic peaks deviate from the accurate positions of the binaries. Consequently, in the iterative-subtraction method, incorrect waveforms of signals $A$ and $B$ would be subtracted, and the maximal $\mathcal{F}$-statistic peak located at a wrong parameter position for overlapped signals is referred to as “the overlapped degeneracy noise”.
@@
Figure 7: The $\mathcal{F}$-statistic degeneracy when two signals overlap. The frequency interval is $5\times\mathrm{d}f$ for Fig. 7(a) and Fig. 7(b), and $4\times\mathrm{d}f$ for Fig. 7(c) and Fig. 7(d). The red crosses marked the real positions of the two binaries.
@@
When the two signals possess different SNRs, the challenge posed by overlapping signals is mitigated. In Fig. 8, we adjust the amplitude $\mathcal{A}$ of binary $B$ to $70\%$ of the value used in Fig. 7(c), resulting in a proportionate decrease in SNR, and the $\mathcal{F}$-statistic value is roughly halved. In this scenario, even though the overlapped signals continue to influence their $\mathcal{F}$-statistic at the two binary positions, the $\mathcal{F}$-statistic value of the two maximal degeneracy noise does not surpass that of binary $A$. Consequently, binary $A$ maintains the maximal $\mathcal{F}$-statistic value in the parameter space. This situation proves advantageous for both our method and the traditional iterative-subtraction method.
@@
We generate two new waveforms using the two wrong maximal $\mathcal{F}$-statistic peaks in Fig. 7(c). The $\mathcal{F}$-statistic degeneracy (Fig. 9) differs from that of binary $A$ and $B$ in Fig. 7(c). Therefore, there is no concern that the real binaries in Fig. 7(c) will be inadvertently removed during the process of addressing degeneracy noise, as detailed in Sec. V.
Figure 9: We generate two signals at the incorrect maximal positions of the $\mathcal{F}$-statistic values in Fig. 7(c) and subsequently calculate the $\mathcal{F}$-statistic degeneracy in the parameter space. Since the $\mathcal{F}$-statistic does not exhibit a high value at the positions of the two binaries in Fig. 7(c), we are not concerned that, in this scenario, removing the individual degeneracy noise in Sec. V.1 will inadvertently subtract the real signal from Fig. 7(c).
III.3.3 Three overlapped signals
@@
It is conceivable for three or more signals with comparable SNR to overlap, especially in scenarios with low SNR and an increased number density of signals. Based on binary $A$, we generate two new binaries—one with a lower frequency (binary $C$) and one with a higher frequency (binary $B$). In Fig. 10(a) and Fig. 10(b), the frequency difference of the three signals is $5\times\mathrm{d}f$, and the interference between these signals is negligible. We anticipate that separating these three signals during the search should be relatively straightforward. However, for the case where the frequency difference of these three signals is $4\times\mathrm{d}f$ (as shown in Fig. 10(c) and Fig. 10(d)), signal interference extinguishes the $\mathcal{F}$-statistic value of binary $A$ and reduces the values of $B$ and $C$. The overlapped signals also generate two erroneously prominent $\mathcal{F}$-statistic peaks above and below the true position of binary $A$ (in Fig. 10(c)). This complicates the search for these three signals, and it is only possible to rediscover binary $A$ after successfully disentangling binaries $B$ and $C$ from the data. How to search out these three overlapped signals is beyond the scope of this article.
@@
Figure 10: The $\mathcal{F}$-statistic degeneracy when three signals overlap. The frequency intervals are $5\times\mathrm{d}f$ for Fig. 10(a) and Fig. 10(b), and $4\times\mathrm{d}f$ for Fig. 10(c) and Fig. 10(d). The red crosses marked the positions of the three binaries.

### v3

@@ III.2 The frequency-position parameter degeneracy
represents the velocity of LISA relative to binary $A$, while $\lambda_{0}$ denotes LISA’s Ecliptic Longitude at the given time. Additionally, $\lambda_{A}$ represents the Ecliptic Longitude of binary $A$, and $\beta_{A}$ represents its Ecliptic Latitude.
Assuming we use binary $A^{\prime}$ as the match waveform, due to the Doppler effect of LISA orbit motion, $f_{A^{\prime}_{\text{do}}}$ is its LISA response frequency
$f_{A^{\prime}_{\text{do}}}=f_{A^{\prime}}\frac{c+\nu_{A^{\prime}}}{c},$ (12)
@@
$\lambda_{A}^{\prime}$ and $\beta_{A}^{\prime}$ are the Ecliptic Longitude and Ecliptic Latitude of binary $A^{\prime}$, respectively. Making these two modulated frequencies equal ($f_{A_{\text{do}}}=f_{A^{\prime}_{\text{do}}}$), we can get the equation
@@
This equation pertains to the presence of a GW signal from binary $A$ in LISA data. When employing the $\mathcal{F}$-statistic to detect the signal, it describes the distribution of the $\mathcal{F}$-statistic in the parameter space.
Considering the aforementioned equation, if the LISA data contains a signal with the frequency $f_{A}$ relative to the SSB, the maximum frequency $f_{A^{\prime}_{\text{max}}}$ in the parameter degeneracy is
$f_{A^{\prime}_{\text{max}}}=f_{A}\frac{c+|\vec{\nu}|}{c-|\vec{\nu}|}\approx 1.0002\times f_{A}$ (16)
@@
when $\theta=\frac{\pi}{2}$, $\lambda_{A^{\prime}}-\lambda_{A}=\pi$, and $\beta_{A}=\beta_{A^{\prime}}=0$. Similarly, the minimum frequency $f_{A^{\prime}_{\text{min}}}$ in the parameter degeneracy is
$f_{A^{\prime}_{\text{min}}}=f_{A}\frac{c-|\vec{\nu}|}{c+|\vec{\nu}|}\approx 0.9998\times f_{A}$ (17)
@@
when $\theta=-\frac{\pi}{2}$, $\lambda_{A^{\prime}}-\lambda_{A}=\pi$, and $\beta_{A}=\beta_{A^{\prime}}=0$. It determines the frequency range $[f_{A^{\prime}_{\text{min}}},\leavevmode\nobreak\ f_{A^{\prime}_{\text{max}}}]$ of the parameter degeneracy generated by the GW signal which has frequency $f_{A}$ relative to the SSB. Eq. 16 and 17 show the frequency range of the parameter degeneracy generated by the $f_{A}$ signal is proportional to $f_{A}$ itself. This results in higher $f_{A}$ exhibiting a more extensive distribution of the $\mathcal{F}$-statistic, while lower $f_{A}$ presents a narrower frequency distribution for the $\mathcal{F}$-statistic.
We selected a set of parameters from the LDC1-4 injection sources catalog and assigned them to binary $A$, with the specific parameter values detailed in Tab. 3 (Appendix A). The parameter degeneracy originating from binary $A$ is plotted in Fig. 4. In Fig. 4(a), $\beta_{A^{\prime}}=\beta_{A}$, and in Fig. 4(b), $\lambda_{A^{\prime}}=\lambda_{A}$. We utilized a spatial resolution of $\pi/40$ radians in the figures. $\theta$ represents the angle between LISA and binary $A$, providing a representation of LISA at various positions in its orbit. The parameter degeneracy for binary $A$ spans a frequency range approximately from $0.9998\times f_{A}$ to $1.0002\times f_{A}$, corresponding to [$2.090406\times 10^{-3}$ Hz, $2.091243\times 10^{-3}$ Hz].
@@
Figure 4: The parameter degeneracy from binary $A$ is derived according to Eq. 15. The parameters of binary $A$$f_{A},\leavevmode\nobreak\ \beta_{A},\leavevmode\nobreak\ \lambda_{A}$—are known, and the position of binary $A$ is denoted by red cross marks. In Fig. 4(a), $\beta_{A}^{\prime}=\beta_{A}$, and in Fig. 4(b), $\lambda_{A}^{\prime}=\lambda_{A}$. The horizontal axis represents the frequency ($f_{A^{\prime}}$) of the matched waveform, binary $A^{\prime}$, while the vertical axis represents its Ecliptic Longitude or Ecliptic Latitude ($\lambda_{A}^{\prime}$ or $\beta_{A}^{\prime}$). $\theta$ denotes the angle between LISA and binary $A$, with different $\theta$ values representing various positions of LISA relative to binary $A$.
@@ III.3.1 Individual signal
We generate a singular LISA response GW signal with the parameters of binary $A$ outlined in Tab. 3. Subsequently, we introduce instrument noise to acquire mock data and compute the parameter degeneracy for data instances with diverse parameter positions. The intentional selection of binary $A$, distinguished by a high SNR, aims to minimize the influence of instrument noise. In Fig. 5, the distribution of degeneracy noise of binary $A$ across three parameter space ($f$, $\beta$, and $\lambda$) is illustrated. Each subfigure has a spatial resolution of $\pi/40$ radians. The colorbar indicates the $\mathcal{F}$-statistic value calculated at the specific parameter position. The red cross in the (2, 5) subfigure marks the real parameter position of binary $A$. The horizontal coordinate represents Ecliptic Longitude $\lambda$, while the vertical coordinate signifies Ecliptic Latitude $\beta$. The frequency interval between two adjacent subfigures corresponds to the minimum frequency resolution $\mathrm{d}f=1.59\times 10^{-8}$ Hz for the two-year detection. From the figure, it is evident that, in addition to the “primary peak” of the $\mathcal{F}$-statistic at the position of binary $A$, numerous “secondary peaks” with elevated $\mathcal{F}$-statistic values emerge. These secondary peaks represent the degeneracy noise mentioned earlier.
@@
Fig. 6(a) and Fig. 6(b) illustrate the distribution of the $\mathcal{F}$-statistic in two-dimensional parameter space. The frequency resolution is $\mathrm{d}f$, and the Ecliptic Latitude $\beta$ or the Ecliptic Longitude $\lambda$ resolution is $\pi/40$ radians. Parameters not displayed in the figures remain constant and correspond to binary $A$. The parameter degeneracy in both figures (Fig. 6(a) and Fig. 6(b)) demonstrates excellent consistency with the theoretical predictions illustrated in Fig. 4(a) and Fig. 4(b).
@@
Figure 6: The red cross in the figures represent the position of binary $A$, while the colorbar indicate the $\mathcal{F}$-statistic value. The vertical coordinate corresponds to Ecliptic Longitude $\lambda$ or Ecliptic Latitude $\beta$, respectively. The parameter degeneracy for binary $A$ spans a frequency range approximately from $0.9998\times f_{A}$ to $1.0002\times f_{A}$, corresponding to [$2.090406\times 10^{-3}$ Hz, $2.091243\times 10^{-3}$ Hz].
III.3.2 Two overlapping signals
@@
Signals that overlap in parameter space introduce complexity to the distribution of the $\mathcal{F}$-statistic. The proximity of two signals in parameter space can lead to interference, manifesting as a high $\mathcal{F}$-statistic peak at an incorrect parameter position, diminishing the value at the actual binary’s parameter position. In scenarios with substantial differences in SNR between the two binaries, the impact of the low SNR signal on the high SNR signal might not be conspicuous, and the low SNR signal could be discerned by subtracting the high SNR signal. However, when overlapping signals exhibit similar SNR, their interference may yield a more pronounced $\mathcal{F}$-statistic peak at an incorrect parameter position. This complicates the iterative-subtraction strategy, as subtracting an incorrect waveform becomes a possibility. In our method, as long as the $\mathcal{F}$-statistic peak at the real binary’s parameter position does not disappear or deviate significantly, the removal of degeneracy noise, detailed in Sec. V, can effectively eliminate the erroneous $\mathcal{F}$-statistic peak.
Building upon binary $A$, we increment its frequency by integer multiples of the minimum frequency resolution $\mathrm{d}f$ to create binary $B$ ($f_{B}=f_{A}+\Delta f=f_{A}+N\times\mathrm{d}f,\leavevmode\nobreak\ N\in Z$). Fig. 7 illustrates the distribution of the $\mathcal{F}$-statistic when two signals coexist. The frequency of binary $B$ in Fig. 7(a) and Fig. 7(b) is increased $5\times\mathrm{d}f$ relative to binary $A$, with red crosses denoting the accurate positions of both binaries—binary $A$ on the left and binary $B$ on the right. When the frequency difference $\Delta f$ between two signals is $\geq 5\times\mathrm{d}f$, the “primary peaks” of the $\mathcal{F}$-statistic for these two signals do not significantly influence each other, each behaving akin to a single source. However, for the binary $B$ in Fig. 7(c) and Fig. 7(d), where the frequency of binary $B$ is incremented by $4\times\mathrm{d}f$ relative to binary $A$, the overlap of signals causes a noticeable deviation in their $\mathcal{F}$-statistic behavior compared to the single signal scenario. In Fig. 7(c), two maximal $\mathcal{F}$-statistic peaks emerge outside the parameter positions of binaries $A$ and $B$, substantially reducing the $\mathcal{F}$-statistic values at the true binary positions. Similarly, in Fig. 7(d), the maximal $\mathcal{F}$-statistic peaks deviate from the accurate positions of the binaries. Consequently, in the iterative-subtraction method, incorrect waveforms of signals $A$ and $B$ would be subtracted, and the maximal $\mathcal{F}$-statistic peak located at a wrong parameter position for overlapping signals is referred to as “the overlapping degeneracy noise”.
@@
Figure 7: The parameter degeneracy when two signals overlap. The frequency interval is $5\times\mathrm{d}f$ for Fig. 7(a) and Fig. 7(b), and $4\times\mathrm{d}f$ for Fig. 7(c) and Fig. 7(d). The red crosses marked the real positions of the two binaries.
@@
When the two signals possess different SNR, the challenge posed by overlapping signals is mitigated. In Fig. 8, we adjust the amplitude $\mathcal{A}$ of binary $B$ to $70\%$ of the value used in Fig. 7(c), resulting in a proportionate decrease in SNR, and the $\mathcal{F}$-statistic value is roughly halved. In this scenario, even though the overlapping signals continue to influence their $\mathcal{F}$-statistic at the two binary positions, the $\mathcal{F}$-statistic value of the two maximal degeneracy noise does not surpass that of binary $A$. Consequently, binary $A$ maintains the maximal $\mathcal{F}$-statistic value in the parameter space. This situation proves advantageous for both our method and the traditional iterative-subtraction method.
@@
III.3.3 Three overlapping signals
@@
It is conceivable for three or more signals with comparable SNR to overlap, especially in scenarios with low SNR and an increased number density of signals. Based on binary $A$, we generate two new binaries—one with a lower frequency (binary $C$) and one with a higher frequency (binary $B$). In Fig. 9(a) and Fig. 9(b), the frequency difference of the three signals is $5\times\mathrm{d}f$, and the interference between these signals is negligible. We anticipate that separating these three signals during the search should be relatively straightforward. However, for the case where the frequency difference of these three signals is $4\times\mathrm{d}f$ (as shown in Fig. 9(c) and Fig. 9(d)), signal interference extinguishes the $\mathcal{F}$-statistic value of binary $A$ and reduces the values of $B$ and $C$. The overlapping signals also generate two erroneously prominent $\mathcal{F}$-statistic peaks above and below the true position of binary $A$ (in Fig. 9(c)). This complicates the search for these three signals, and it is only possible to rediscover binary $A$ after successfully disentangling binaries $B$ and $C$ from the data. How to search out these three overlapping signals is beyond the scope of this article.
@@
Figure 9: The parameter degeneracy when three signals overlap. The frequency intervals are $5\times\mathrm{d}f$ for Fig. 9(a) and Fig. 9(b), and $4\times\mathrm{d}f$ for Fig. 9(c) and Fig. 9(d). The red crosses marked the positions of the three binaries.
III.4 The overlapping signals from LDC1-4 injection sources
We identified injection sources with SNR $\geq 7$ in LDC1-4 data and evaluated the presence of overlapping sources among them for two-year detection. These injection sources were categorized based on SNR thresholds: SNR $\geq 15$ or SNR ranging from $7$ to $15$. In Tab. 1, “Sole” indicates that no other source occurs within its own $5\times\mathrm{d}f$ range, thus not considered an overlapping source. “Two” indicates a source overlapping within the $5\times\mathrm{d}f$ range, as illustrated in Fig. 7(c) and Fig. 7(d). “Three” denotes three continuous overlapping sources, where the frequency interval between two sources is less than $5\times\mathrm{d}f$, and the interval between one of the sources and the third source is also less than $5\times\mathrm{d}f$, as depicted in Fig. 9(c) and Fig. 9(d). “Four” indicates four continuous overlapping sources in a sequence, and so on. Among the injection sources with SNR $\geq 15$, we found a maximum of five continuous overlapping sources, with non-overlapping sources comprising 73.6%. For injection sources with SNR between 7 and 15, up to ten continuous overlapping sources were observed, with non-overlapping sources accounting for only 47.2%. It is evident that as SNR decreases, the likelihood of overlapping sources increases, impacting our signal processing and reducing the detection rate of the search pipeline. In fact, we expect as the detection time increases, more and more signals can be resolved, which means that there will be more and more overlap problems.
Injection sources’ SNR Sole Two Three Four Five Six Seven Eight Nine Ten Total number
$\geq 15$ 8079 2208 561 124 10 0 0 0 0 0 10982
$7\sim 15$ 7874 4344 2175 1136 560 318 154 72 45 20 16698
Table 1: The number of overlapping signals from LDC1-4 injection sources is categorized as follows: “Sole” denotes no overlapping source. “Two” indicates that the frequency interval between two sources is less than $5\times\mathrm{d}f$, resulting in their overlap. “Three” implies that not only the frequency interval between two sources is less than $5\times\mathrm{d}f$, but also the frequency interval between one of the sources and the third source is less than $5\times\mathrm{d}f$, and so forth.